OpenStudy (anonymous):

Simplify the expression. –20 • 1/4 • 8 + 12 ÷ 4

OpenStudy (notamathgenius):

Do you know how to do pemdas?

OpenStudy (anonymous):

-20*2+3=-40+3=-37